(1) Write the net ionic equation for the reaction that occurs when equal volumes of 0.191 M aqueous hypochlorous acid and ammonia are mixed. It is not necessary to include states such as (aq) or (s).

At equilibrium will the reactants or products be favored.

The pH of the resulting solution will be greater,less, or equal to 7.
Answer #1

hypochlorous (HOCl) is weak acid acid and ammonia (NH3) is weak base.

Net ionic equation is

HOCl (aq) + NH3 (aq) <........................> NH4+ (aq) + OCl- (aq)

here

As both HOCl and NH3 are weak. so At equilibrium will the reactants be favored.

pKA (HOCl) = 7.456

pKb of NH3 = 4.75

[NH4OCl] = 0.191 / 2 = 0.0955 M

pH = 0.5 * (pKw + pKa - pkb) = 0.5 (14 + 7.456 - 4.725) = 8.37 > 7.0

thus

pH of the resulting solution will be greater than 7.